# AppReferences

### v2

@@ 3.1 Causal Estimands
To define our causal estimands of interest, we postulate that the units can be marginally grouped into clusters with finite first moments with respect to the treatment level. Consequently, each unit $i$ can belong to: (i) the dissociative stratum if $S_{i}^{(1)}=S_{i}^{(0)}$, which is equivalent to $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]=\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$; or (ii) the associative positive stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]>\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (throughout the paper, we indicate this condition as $S_{i}^{(1)}\succ S_{i}^{(0)}$); or (iii) the associative negative stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]<\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (throughout, we indicate this condition as $S_{i}^{(1)}\prec S_{i}^{(0)}$).
Definition 3.1 (Causal Estimands).
Conditionally on the strata allocation, we define the causal estimands for the expected dissociative effect (EDE) and the expected associative effects (EAE) as
$EDE=\tau_{0}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}=S_{i}^{(0)}],$
$EAE_{+}=\tau_{+1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\succ S_{i}^{(0)}],$
$EAE_{-}=\tau_{-1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\prec S_{i}^{(0)}].$
@@
Theorem 3.2.
Invoking assumptions (2.1) and (2.1), the expected value of each potential outcome can be rewritten as the following
$\mathbb{E}\big{[}Y_{i}(t)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{]}$
$\quad=\int_{p_{0}p_{1}}\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$
×$\quad\quad\times Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{)}dp_{0}p_{1}dx;$ (8)
where the inner expectations $\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{)}$, that is defined by the model for the potential post-treatment variables.
@@ 3.2 Discovery of Principal Strata
While Bayesian nonparametric mixtures excel at capturing cluster structures, our focus shifts to leveraging this framework for causal inference. To estimate our causal estimands of interest, we need to characterize the joint distribution of $\{S_{i}^{(0)},S_{i}^{(1)}\}$ so that we can determine each unit’s allocation to a principal strata. This joint distribution cannot be directly observed due to the fundamental problem of causal inference. Nevertheless, our proposed Bayesian nonparametric model allows for robust estimation of this joint distribution through its hierarchical construction and shared parameters, effectively addressing this problem. In this section, we illustrate how our proposed Bayesian nonparametric model enables the characterization of principal strata in a statistically robust and fully data-adaptive manner.
@@
Our proposed model is defined such that the atoms $\{\psi_{l}\}_{l\geq 1}$ are shared between the two potential post-treatment outcomes for the same unit. This formulation yields precise posterior probabilities that quantify, for each unit $i$, whether the latent indicators $\{S_{i}^{(0)},S_{i}^{(1)}\}$ map to identical or distinct mixture components. Specifically, for any unit $i\in\{1,\ldots,n\}$, the probability of membership in the dissociative stratum—where $S_{i}^{(0)}\neq S_{i}^{(1)}$—equals the probability that the atom under control differs from the atom under treatment. This probability is provably non-zero and can be derived a priori through the following theorem.
Theorem 3.3.
@@
$\Pr[S_{i}^{(0)}=S_{i}^{(1)}]=\frac{\rho_{2}(x_{i})[(1+\rho_{2}(x_{i})-2\rho_{1}(x_{i}))^{L}-1]}{\rho_{2}(x_{i})-2\rho_{1}(x_{i})}$ (9)
where $\rho_{1}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}\big{]}$ and $\rho_{2}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}^{2}\big{]}$ and $0\leq\Pr[S_{i}^{(0)}=S_{i}^{(1)}]\leq 1$.
@@
[…]
@@ 3.3 Posterior Inference
Theorem 3.2 establishes that principal causal estimands are characterized as functions of both the primary response and potential post-treatment outcomes, each involving specific model parameters. Following the Bayesian paradigm, we estimate these parameters through their posterior distributions conditional on the observed data. This approach provides not only point estimates but also complete uncertainty quantification.
Sampling from the posterior joint distribution is straightforward via Gibbs sampling. In particular, the proposed algorithm (described in detail in the Supplementary Material) takes inspiration from Teh et al. (2004, 2006); Teh & Jordan (2009) for the hierarchical structure, and from Fasano et al. (2022) for the probit regression in the weights. We assume a general linear regression for the outcome model that depends on the parameter $\theta_{0}^{Y}$ for the outcome under control and $\theta_{1}^{Y}$ for the treated outcome.
@@
    $\Pr\{S_{i}^{(t)}=l\}\propto\pi_{l}^{(t)}(\mathbf{x}_{i})\Pr(P_{i}(t)\mid S_{i}^{(t)}=l,X_{i};\eta_{l},\sigma_{l}),$
    where $\pi_{l}^{(t)}(\mathbf{x}_{i})=\Phi(\alpha_{l}^{(t)}(\mathbf{x}_{i}))\prod_{r<l}(1-\Phi(\alpha_{r}^{(t)}(\mathbf{x}_{i})))p$ for $l=1,\dots,L-1$ and $\Phi(\alpha_{L}^{(t)}(\mathbf{x}_{i}))=1$.
@@
    $\psi_{l}\propto\Pi_{x}\prod_{i:S_{i}^{(0)}=l}\Pr(P_{i}(0)\mid S_{i}^{(0)}=l,X_{i};\psi_{l})\prod_{i:S_{i}^{(1)}=l}\Pr(P_{i}(1)\mid S_{i}^{(1)}=l,X_{i};\psi_{l}).$
@@
    $\beta^{(t)}\propto\Pr(\beta^{(t)})\prod_{i=1}^{n}\Pr(P_{i}(t)\mid S_{i}^{(t)},X_{i};\beta^{(t)})$
    following (Fasano et al., 2022) demonstration, detail in the Supplementary Material.
@@
    $P_{i}(1-T_{i})\propto\Pr(Y_{i}\mid P_{i}(0),P_{i}(1),X_{i};\theta_{0}^{Y})^{1-T_{i}}\Pr(Y_{i}\mid P_{i}(0),P_{i}(1),X_{i};\theta_{1}^{Y})^{T_{i}}\Pr(P_{i}(1-T_{i})\mid X_{i};\psi).$
  • The posterior distribution for the outcome model is built from
    $\theta_{t}^{Y}\sim\Pr(\theta_{t}^{Y})\prod_{i=1}^{n}\text{P}(Y_{i}(t)\mid P_{i}(0),P_{i}(1),X_{i};\theta_{t}^{Y}).$
@@ 4 Simulation Study
The performance of the proposed mixture model is evaluated through a simulation study. Our objective is to investigate the model’s ability to (i) accurately impute the missing post-treatment and outcome variables, that is, to assess the potential bias of the expected value $E[P_{i}(1)-P_{i}(0)]$ and $E[Y_{i}(1)-Y_{i}(0)]$ over the units $i=1\dots,n$, (ii) correctly identify the principal strata and estimate the principal causal effects. To achieve this, we conducted simulations under five different data-generating mechanisms and analyzed the results to understand the model’s behavior in different scenarios.
@@
$\begin{bmatrix}Y(0)\\ Y(1)\end{bmatrix}\sim\mathcal{N}_{2}\begin{pmatrix}\begin{bmatrix}\theta_{00}+\theta_{01}P(0)\\ \theta_{10}+\theta_{11}P(1)+\theta_{12}P(0)+\theta_{13}P(0)P(1)\end{bmatrix},\begin{bmatrix}e^{\lambda_{0}}&0\\ 0&e^{\lambda_{0}+\lambda_{1}P(1)}\end{bmatrix}\end{pmatrix}.$ (10)
@@
[…]
with $\mu_{\theta}=0$, $\sigma_{\theta}=10$, $\mu_{\lambda}=0$, and $\sigma_{\lambda}=2$. Clearly, our model can accommodate more complex choices for the outcome, but we decided to opt for a simpler model following what had already been done in the literature by Schwartz et al. (2011) and Lu et al. (2023).
For the data generating process, we simulate two Bernoulli confounders $(X_{1},X_{2})$ for scenarios 1 to 4 and five Bernoulli confounders $(X_{1},X_{2},X_{3},X_{4},X_{5})$ for scenario 5, and a binary treatment variable such that $T_{i}\sim\mbox{Be}(\mbox{logit}(0.4X_{1i}+0.6X_{2i}))$ for the scenarios $1-4$ and $T_{i}\sim\mbox{Be}(\mbox{logit}(0.4X_{1i}+0.6X_{2i}-0.3X_{3i}+0.2X_{4i}X_{5i}))$ for scenario 5, for $i=1,\dots,n$.
@@
$P_{i}(0)|S_{i}^{(0)}=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}),\quad P_{i}(1)|S_{i}^{(1)}=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}).$
@@
$\beta^{(t)}\sim\mathcal{N}_{(p+1)(L-1)}(0,20\times\mathbf{1}_{(p+1)(L-1)}),$
@@
for $t\in\{0,1\}$, $l\in\{1,\dots,20\}$, and $p$ according with the covariates considered in different settings, and where $\mathbf{1}_{q}$ is a diagonal matrix $q\times q$.
@@
The results for the five scenarios demonstrate the strong ability of our proposed model to impute missing variables and accurately capture the true distribution across different data generating processes. The medians of the bias are close to zero and the interquartile range is reasonable given the simulated variability. In contrast, the two competing models highlight their limitations in recovering the same information. The algorithm from Lu et al. (2023) does not estimate the distribution of the potential outcome for the post-treatment variable, while the model from Schwartz et al. (2011) exhibits significant bias in four out of five scenarios and a higher interquartile range than our model across all scenarios. Furthermore, the comparison of $\mathbb{E}[Y_{i}(1)-Y_{i}(0)]$ reveals substantial bias in all five scenarios for both competing models Lu et al. (2023) and Schwartz et al. (2011).
[…]
@@
Bias comparison of the two methods based on different simulation scenarios. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 $\mathbb{E}[P(1)-P(0)]$ Our proposed method Median -0.0164 0.0027 -0.0318 0.0027 0.0011 IQR 0.0315 0.0254 0.0462 0.0325 0.0183 Schwartz et al. (2011) Median -0.0637 0.3388 0.1254 0.3280 0.3572 IQR 0.0989 0.0986 0.0956 0.0963 0.1216 Lu et al. (2024) Median IQR $\mathbb{E}[Y(1)-Y(0)]$ Our proposed method Median 0.0024 0.0256 0.0067 0.0419 0.0344 IQR 0.2235 0.1550 0.4661 0.1357 0.1316 Schwartz et al. (2011) Median -1.9891 -1.8265 -1.4363 -1.3115 -1.8856 IQR 0.2546 0.2110 0.1904 0.1836 0.3727 Lu et al. (2024) Median -1.3491 0.3070 -0.5848 0.1393 0.0751 IQR 0.1092 0.1206 0.1009 0.0869 0.1078 {tabnote} IQR: interquartile range.
@@
To assess the accurate identification of the principal strata, we use the adjusted Rand index (Hubert & Arabie, 1985) and evaluate the estimation of the principal causal effects. The adjusted Rand index values for the five scenarios are reported in Table 4. For all scenarios, the index is close to $1$ and confirms that our proposed model can correctly identify the principal strata and, combined with the good missing data imputation, allows us to estimate the expected associative and dissociative effects. Table 4 shows that our model correctly identifies the number of principal strata—two in Scenario 1 and 3, three in the others—while also providing accurate causal effect estimates. In contrast, neither the model by Lu et al. (2023) nor that of Schwartz et al. (2011)identify the principal strata according to our definition.
[…]
@@
Adjusted rand index for the five simulated scenarios computed on the point estimated partitions obtained with our proposed model. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Mean 0.9850 0.9906 0.9706 0.9717 0.9154 Standard deviation 0.0997 0.0597 0.1021 0.0892 0.1577
[…]
@@
True and estimated principal causal effects—mean and standard deviation in the brackets—from our proposed method across scenarios. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 True Estimated True Estimated True Estimated True Estimated True Estimated $\tau_{-}$ NA $-1.00$ $-0.98$ (0.32) NA $1.50$ $1.46$ (0.33) $-1.00$ $-0.79$ (0.68) $\tau_{0}$ $0.50$ $0.11$ (0.56) $4.00$ $4.03$ (0.61) $5.60$ $4.98$ (0.52) $4.00$ $4.11$ (0.53) $4.00$ $4.57$ (1.50) $\tau_{+}$ $9.00$ $8.79$ (1.25) $9.00$ $9.02$ (0.45) $8.60$ $9.36$ (0.87) $6.50$ $6.59$ (0.35) $9.00$ $8.91$ (0.44) {tabnote} NA: Not applicable. The stratum was not present in the scenario.

### v3

@@ 3.2 Causal Estimands
To define our causal estimands of interest, we postulate that the units can be marginally grouped into clusters with finite first moments with respect to the treatment level. Consequently, each unit $i$ can belong to: (i) the dissociative stratum if $S_{i}^{(1)}=S_{i}^{(0)}$, which is equivalent to $\mathbb{E}_{S_{i}^{(1)}}\{P_{i}(1)\}=\mathbb{E}_{S_{i}^{(0)}}\{P_{i}(0)\}$; or (ii) the associative positive stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}\{P_{i}(1)\}>\mathbb{E}_{S_{i}^{(0)}}\{P_{i}(0)\}$ (throughout the paper, we indicate this condition as $S_{i}^{(1)}\succ S_{i}^{(0)}$); or (iii) the associative negative stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}\{P_{i}(1)\}<\mathbb{E}_{S_{i}^{(0)}}\{P_{i}(0)\}$ (throughout, we indicate this condition as $S_{i}^{(1)}\prec S_{i}^{(0)}$).
Definition 1 (Causal Estimands)
Conditionally on the strata allocation, we define the causal estimands for the expected dissociative effect and the expected associative effects as
$\tau_{0}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}=S_{i}^{(0)}\},$
$\tau_{+}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\succ S_{i}^{(0)}\},$
$\tau_{-}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\prec S_{i}^{(0)}\}.$
@@
Theorem 1
Invoking assumptions (1) and (2), the expected value of each potential outcome can be rewritten as the following
$\mathbb{E}\big\{Y_{i}(t)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big\}$
$\quad=\int_{p_{0}p_{1}}\mathbb{E}\big\{Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big\}$
$\quad\quad\times\text{P}\big(P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big)dp_{0}p_{1}dx;$ (8)
where the inner expectations $\mathbb{E}\big\{Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big\}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\mbox{pr}\big(P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big)$, that is defined by the model for the potential post-treatment variables.
@@ 3.3 Discovery of Principal Strata
While Bayesian nonparametric mixtures excel at capturing cluster structures, our focus shifts to leveraging this framework for causal inference. To estimate our causal estimands of interest, we need to characterize the joint distribution of $\{S_{i}^{(0)},S_{i}^{(1)}\}$ so that we can determine each unit’s allocation to a principal strata. This joint distribution cannot be directly observed due to the fundamental problem of causal inference. Nevertheless, our proposed Bayesian nonparametric model allows for robust estimation of this joint distribution through its hierarchical construction and shared parameters, effectively addressing this problem. In this section, we illustrate how CASBAH enables the characterization of principal strata in a statistically robust and fully data-adaptive manner.
@@
CASBAH is defined such that the atoms $\{\psi_{l}\}_{l\geq 1}$ are shared between the two potential post-treatment outcomes for the same unit. This formulation yields precise posterior probabilities that quantify, for each unit $i$, whether the latent indicators $\{S_{i}^{(0)},S_{i}^{(1)}\}$ map to identical or distinct mixture components. Specifically, for any unit $i\in\{1,\ldots,n\}$, the probability of membership in the dissociative stratum—where $S_{i}^{(0)}\neq S_{i}^{(1)}$—equals the probability that the atom under control differs from the atom under treatment. This probability is provably non-zero and can be derived a priori through the following theorem.
Theorem 2
@@
$\text{P}\{S_{i}^{(0)}=S_{i}^{(1)}\}=\frac{\rho_{2}(x_{i})[\{1+\rho_{2}(x_{i})-2\rho_{1}(x_{i})\}^{L}-1]}{\rho_{2}(x_{i})-2\rho_{1}(x_{i})}$ (9)
where $\rho_{1}(x_{i})=\mathbb{E}\big[\Phi\big(\alpha(x_{i})\big)\big]$ and $\rho_{2}(x_{i})=\mathbb{E}\big[\Phi\big(\alpha(x_{i})\big)^{2}\big]$ and $0\leq\text{P}\{S_{i}^{(0)}=S_{i}^{(1)}\}\leq 1$.
@@
Such probability varies according to the function $\alpha(x)=\beta_{0}+\beta_{1}x$. Thus, the probability of belonging to the dissociative stratum depends on the observed covariates $x_{i}$ for each unit $i$ and also on the posterior distribution of the parameters $\beta=\{\beta_{0},\beta_{1}\}$. The $\beta$ posterior distributions are a function of all observed data, such that the probability of belonging to the dissociative strata is directly characterized from the data.
@@ 3.4 Posterior Inference
Theorem 1 establishes that principal causal estimands are characterized as functions of both the primary response and potential post-treatment outcomes, each involving specific model parameters. Following the Bayesian paradigm, we estimate these parameters through their posterior distributions conditional on the observed data. This approach provides not only point estimates but also complete uncertainty quantification.
Sampling from the posterior joint distribution is straightforward via Gibbs sampling. In particular, the proposed algorithm (described in detail in the Supplementary Material) takes inspiration from Teh et al. (2004, 2006); Teh and Jordan (2009) for the hierarchical structure, and from Fasano et al. (2022) for the probit regression in the weights. We assume a general linear regression for the outcome model that depends on the parameter $\theta_{0}^{Y}$ for the outcome under control and $\theta_{1}^{Y}$ for the treated outcome.
@@
    $\text{P}\{S_{i}^{(t)}=l\}\propto\pi_{l}^{(t)}(\mathbf{x}_{i})\text{P}(P_{i}(t)\mid S_{i}^{(t)}=l;\eta_{l},\sigma_{l}),$
    where $\pi_{l}^{(t)}(\mathbf{x}_{i})=\Phi(\alpha_{l}^{(t)}(\mathbf{x}_{i}))\prod_{r<l}\{1-\Phi(\alpha_{r}^{(t)}(\mathbf{x}_{i}))\}$ for $l=1,\dots,L-1$ and $\Phi(\alpha_{L}^{(t)}(\mathbf{x}_{i}))=1$.
@@
    $\psi_{l}\propto\Pi_{x}\prod_{i:S_{i}^{(0)}=l}\text{P}(P_{i}(0)\mid S_{i}^{(0)}=l,X_{i};\psi_{l})\prod_{i:S_{i}^{(1)}=l}\text{P}(P_{i}(1)\mid S_{i}^{(1)}=l,X_{i};\psi_{l}).$
@@
    $\beta^{(t)}\propto\text{P}(\beta^{(t)})\prod_{i=1}^{n}\text{P}(P_{i}(t)\mid S_{i}^{(t)},X_{i};\beta^{(t)})$
    following (Fasano et al., 2022). Details are reported in the Supplementary Material.
@@
    $P_{i}(1-T_{i})\propto\text{P}(Y_{i}\mid P_{i}(0),P_{i}(1),X_{i};\theta_{0}^{Y})^{1-T_{i}}\text{P}(Y_{i}\mid P_{i}(0),P_{i}(1),X_{i};\theta_{1}^{Y})^{T_{i}}\text{P}(P_{i}(1-T_{i})\mid X_{i};\psi).$
  • The conditional posterior distribution for the outcome model is built from
    $\theta_{t}^{Y}\sim\text{P}(\theta_{t}^{Y})\prod_{i=1}^{n}\text{P}(Y_{i}(t)\mid P_{i}(0),P_{i}(1),X_{i};\theta_{t}^{Y}).$
@@ 4 Simulation Study
The performance of the proposed mixture model is evaluated through a simulation study. Our objective is to investigate the model’s ability to (i) accurately impute the missing post-treatment and outcome variables, that is, to assess the potential bias of the expected value $E\{P_{i}(1)-P_{i}(0)\}$ and $E\{Y_{i}(1)-Y_{i}(0)\}$ over the units $i=1\dots,n$, (ii) correctly identify the principal strata and estimate the principal causal effects. To achieve this, we conducted simulations under five different data-generating mechanisms and analyzed the results to understand the model’s behavior in different scenarios.
@@
$\begin{bmatrix}Y_{i}(0)\\ Y_{i}(1)\end{bmatrix}\sim\mathcal{N}_{2}\begin{pmatrix}\begin{bmatrix}\theta_{00}+\theta_{01}P_{i}(0)\\ \theta_{10}+\theta_{11}P_{i}(1)+\theta_{12}P_{i}(0)+\theta_{13}P(0)P_{i}(1)\end{bmatrix},\begin{bmatrix}e^{\lambda_{0}}&0\\ 0&e^{\lambda_{0}+\lambda_{1}P_{i}(1)}\end{bmatrix}\end{pmatrix}$. (10)
@@
$\theta^{(0)}=(\theta_{00},\theta_{01})\sim\mathcal{N}_{2}(\mu_{\theta},\sigma^{2}_{\theta}I_{2})$ $\mbox{ and }\theta^{(1)}=(\theta_{10},\theta_{11},\theta_{12},\theta_{13})\sim\mathcal{N}_{4}(\mu_{\theta},\sigma^{2}_{\theta}I_{4});$
$\lambda_{0}\sim\mathcal{N}(\mu_{\lambda},\sigma^{2}_{\lambda})$ $\mbox{ and }\lambda_{1}\sim\mathcal{N}(\mu_{\lambda},\sigma^{2}_{\lambda});$
with $\mu_{\theta}=0$, $\sigma_{\theta}=10$, $\mu_{\lambda}=0$, and $\sigma_{\lambda}=2$. Clearly, CASBAH can accommodate more complex choices for the outcome, but we decided to opt for a simpler model following what had already been done in the literature by Schwartz et al. (2011) and Lu et al. (2023).
For the data generating process, we simulate two Bernoulli confounders $(X_{1},X_{2})$ for Scenarios 1 to 4 and five Bernoulli confounders $(X_{1},X_{2},X_{3},X_{4},X_{5})$ for Scenario 5, and a binary treatment variable such that $T_{i}\sim\mbox{Be}(\mbox{logit}(0.4X_{1i}+0.6X_{2i}))$ for Scenarios 141-4 and $T_{i}\sim\mbox{Be}(\mbox{logit}(0.4X_{1i}+0.6X_{2i}-0.3X_{3i}+0.2X_{4i}X_{5i}))$ for Scenario 5, for $i=1,\dots,n$.
@@
$\{P_{i}(0)|S_{i}^{(0)}=s\}\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}),\qquad\{P_{i}(1)|S_{i}^{(1)}=s\}\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}).$
@@
$\beta^{(t)}\sim\mathcal{N}_{(p+1)(L-1)}(0,20\times I_{(p+1)(L-1)}),$
@@
for $t\in\{0,1\}$, $l\in\{1,\dots,20\}$, and $p$ according with the covariates considered in different settings, and where $I_{q}$ is a diagonal matrix $q\times q$.
@@
The results for the five scenarios demonstrate the strong ability of CASBAH to impute missing variables and accurately capture the true distribution across different data generating processes. The medians of the bias are close to zero and the interquartile range is reasonable given the simulated variability. In contrast, the two competing models highlight their limitations in recovering the same information. The algorithm from Lu et al. (2023) does not estimate the distribution of the potential outcome for the post-treatment variable, while the model from Schwartz et al. (2011) exhibits significant bias in four out of five scenarios and a higher interquartile range than CASBAH across all scenarios. Furthermore, the comparison of $\mathbb{E}[Y_{i}(1)-Y_{i}(0)]$ reveals substantial bias in all five scenarios for both competing models Lu et al. (2023) and Schwartz et al. (2011).
@@
Table 1: Bias comparison of the three methods based on different simulation scenarios. (IQR: interquartile range.)
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5
$\mathbb{E}\{P(1)-P(0)\}$
CASBAH Median -0.0142 0.0010 -0.0032 0.0002 0.0019
IQR 0.0235 0.0159 0.0285 0.0234 0.0193
Schwartz et al. (2011) Median -0.0637 0.3388 0.1254 0.3280 0.3572
IQR 0.0989 0.0986 0.0956 0.0963 0.1216
Lu et al. (2024) Median
IQR
$\mathbb{E}\{Y(1)-Y(0)\}$
CASBAH Median -0.0534 -0.0001 -0.0159 -0.0027 0.0020
IQR 0.0699 0.1049 0.0885 0.0909 0.0720
Schwartz et al. (2011) Median -1.9891 -1.8265 -1.4363 -1.3115 -1.8856
IQR 0.2546 0.2110 0.1904 0.1836 0.3727
Lu et al. (2024) Median -1.3491 0.3070 -0.5848 0.1393 0.0751
IQR 0.1092 0.1206 0.1009 0.0869 0.1078
@@
To assess the accurate identification of the principal strata, we use the adjusted Rand index (Hubert and Arabie, 1985) and evaluate the estimation of the principal causal effects. The adjusted Rand index values for the five scenarios are reported in Table 2. For all scenarios, the index is close to $1$ and confirms that CASBAH can correctly identify the principal strata and, combined with the good missing data imputation, allows us to estimate the expected associative and dissociative effects. Table 3 shows that CASBAH correctly identifies the number of principal strata—two in Scenario 1 and 3, three in the others—while also providing accurate causal effect estimates. In contrast, neither the model by Lu et al. (2023) nor that of Schwartz et al. (2011)identify the principal strata according to our definition.
@@
Table 2: Adjusted rand index for the five simulated scenarios computed on the point estimated partitions obtained with our proposed model.
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5
Mean 0.9850 0.9906 0.9706 0.9717 0.9154
Standard deviation 0.0997 0.0597 0.1021 0.0892 0.1577
@@
Table 3: True and estimated principal causal effects—mean and standard deviation in the brackets—from CASBAH model across scenarios. (NA: Not applicable. The stratum was not present in the scenario.)
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5
True Estimated True Estimated True Estimated True Estimated True Estimated
$\tau_{-}$ N/A 1.00-1.00 0.98-0.98 (0.32) N/A $1.50$ $1.46$ (0.33) 1.00-1.00 0.79-0.79 (0.68)
$\tau_{0}$ 0.5-0.5 0.50-0.50 (0.24) $4.00$ $4.03$ (0.61) $5.60$ $4.98$ (0.52) $4.00$ $4.11$ (0.53) $4.00$ $4.57$ (1.50)
$\tau_{+}$ 33 $2.91$ (0.11) $9.00$ $9.02$ (0.45) $8.60$ $9.36$ (0.87) $6.50$ $6.59$ (0.35) $9.00$ $8.91$ (0.44)